1.Sketch the area under the standard normal curve over the indicated interval and find the specified area. (Round your answer to four decimal places.)
a. The area to the right of z = 1.54 is?
b. The area to the left of z = 0.37 is?
c. The area to the left of z = −1.31 is?
2. Sketch the area under the standard normal curve over the indicated interval and find the specified area. (Round your answer to four decimal places.)
The area between
a. z = −1.50 and z = 1.90 is?
b. z = 0 and z = −1.86 is?

Answer by ewatrrr(24785)


Shade to right of z = 1.54
p(z > 1.54) = 1- p(z <= 1.54) = 1 - normalcdf(-10,1.54) = .0617

Shade to Left of z = .37
p( z < .37) = normalcdf(-10,.37) = .6443

Shade to Left of z = -1.31
p(z < -1.31) = normalcdf(-10,-1.91) = .0951
Since this is a continuous function, we have P(z < value) = P(z ≤ value).
|
2 Sketch the area under the standard normal curve between the z-values given.
P(-1.50 ≤ z ≤ 1.90) = normalcdf(-1.50, 1.90)
P(-1.86 ≤ z ≤ 0) = normalcdf(-1.86, 0)